# The integral Chow ring of $\mathcal{R}_{2}$

### v1

@@ 2.4 Proof of Theorem 2
    for every object $Y\in\mathcal{Y}(S)$ over $S$ there is a covering $\widetilde{S}\to S$ and an object $\widetilde{X}\in\mathcal{X}^{\mathrm{pre}}(\widetilde{S})$ such that $\alpha^{\mathrm{pre}}(\widetilde{X})=Y_{|\widetilde{S}}$,
@@
We start with showing that $\alpha^{\mathrm{pre}}$ is fully faithfull.
@@
The morphism $\alpha^{\mathrm{pre}}$ is fully faithfull.
@@
Since the second map is fully faithfull by construction, we only need to prove that the first map is faithfull. We will refer to again by the symbol $\alpha^{\mathrm{pre}}$. It is enough to prove that $\alpha^{\mathrm{pre}}(S):\mathcal{X}^{\mathrm{pre}}(S)\to\mathcal{R}_{2}^{\mathrm{st}_{1}}(S)$ is fully faithfull for all schemes $S$. Let the $u,u^{\prime}:S\to\mathcal{X}^{\mathrm{pre}}$ objects in $\mathcal{X}$ and let $\alpha^{\mathrm{pre}}(u)=(\pi:C\to S,\eta,\beta)$ and $\alpha^{\mathrm{pre}}(u^{\prime})=(\pi^{\prime}:C^{\prime}\to S,\eta^{\prime},\beta^{\prime})$ be their images under $\alpha^{\mathrm{pre}}$. We wish to show that the induced map
$\mathrm{Mor}_{\mathcal{X}(S)}(u,u^{\prime})\to\mathrm{Mor}_{\mathcal{R}_{2}^{\mathrm{st}_{1}}(S)}((C/S,\eta,\beta),(C^{\prime}/S^{\prime},\eta^{\prime},\beta^{\prime}))$ (19)
@@
[…]
commutes. The uniqueness of $\phi$ is clear. To show the existence of $\phi$, we consider the corresponding diagram where the vertical arrow are the canonical maps and the lower horizontal arrow is given by the pullback of forms via $\varphi^{*}$
${C}$${C^{\prime}}$${\mathbb{P}\,(\pi_{*}\omega_{C/S})^{\vee}}$${\mathbb{P}\,(\pi_{*}\omega_{C^{\prime}/S})^{\vee}}$$\scriptstyle{\varphi}$$\scriptstyle{p}$$\scriptstyle{p^{\prime}}$$\scriptstyle{\varphi^{*}}$
@@
[…]
where $h$ and $h^{\prime}$ are isomorphisms. We have obtained in this way an isomorphism $\widetilde{\phi}=h^{\prime}\circ\varphi^{*}\circ h^{-1}:\mathbb{P}^{1}_{\widetilde{S}}\to\mathbb{P}^{1}_{\widetilde{S}}$ making (20) to commute.
@@
    isomorphism $\varphi$ between the curve $C$ in (9) and $C^{\prime}$;
@@
Let $\bar{s}$ be a geometric point of $S$. We will prove that (a) and (b) holds in an étale neighborhood of $\bar{s}$. By [Vis98], we know that, perhaps after base-change by an étale cover of $S$, there exists $F$ in $\mathcal{X}^{\mathrm{pre}}$ over $S$ such that there is an isomorphism $\varphi:C\to C^{\prime}$ over $S$. By Lemma 3, we have $\eta^{\prime}_{\bar{s}}=q_{\bar{s}}^{\prime*}\mathcal{O}(1)\otimes\mathcal{O}_{C^{\prime}_{\bar{s}}}(-w_{i}-w_{j})$ for exactly one subset $\{w_{i},w_{j}\}$ of cardinality $2$ of the set of Weierstrasse points of $C^{\prime}_{\bar{s}}$. Here $q^{\prime}_{\bar{s}}:C^{\prime}_{\bar{s}}\to\mathbb{P}^{1}$ is any degree $2$ map such that $q_{\bar{s}}^{\prime}(\{w_{i},w_{j}\})=\{0,\infty\}\subseteq\mathbb{P}^{1}$. Up to composing the map (10) with an automorphism of $\mathbb{P}^{1}$ fixing $\{0,\infty\}$ (which amounts to change $F$but not $C$- ), we may assume that
@@
[…]
@@
Theorem 2 now follows immediately from Lemma 23 and Propositions 24 and 25. ∎
@@ 3 Computation of the Chow ring of $\mathcal{R}_{2}$
where $\underline{\Delta}\subseteq\mathbb{P}\,\mathrm{Sym}^{4}(V^{\vee})$ is the locus of polynomials (up to non-zero scalars) having a double root or a root at 00 or $\infty$.
@@
where $h=c_{1}^{G}(\mathcal{O}_{\mathbb{P}\,\mathrm{Sym}^{4}(V^{\vee})}(1))$.
@@
$\mathrm{CH}_{*}^{G}(\Delta)\to\mathrm{CH}_{*}^{G}(\mathbb{P}\,\mathrm{Sym}^{4}(V^{\vee}))\to\mathrm{CH}^{G}_{*}(\mathbb{P}\,\mathrm{Sym}^{4}(V^{\vee})\smallsetminus\underline{\Delta})\to 0$ (22)
@@ 3.1 A $G$-equivariant envelope for $\underline{\Delta}$
    if $x$ is a $k$-valued point of $X_{i}\smallsetminus X_{i+1}$, then there is a unique $k$-valued point $y$ of $X_{i}^{\prime}$ mapping to $x$.
then the map $p=\sqcup p_{i}:\sqcup_{i=1}^{d}X_{i}\to X$ is a $G$-equivariant envelope.
@@
by $(F,G)\mapsto F^{2}G$, where $G\in\{X^{2},Y^{2},XY\}$ We set $p_{3}=p_{3,1}\sqcup p_{3,2}$ and $\underline{\Delta}_{3}\subseteq\underline{\Delta}_{2}$ to be the image of $p_{3}$.
@@
We will check the conditions of Lemma 29. Condition $(i)$ and $(ii)$ are obvious. We will check that condition $(iii)$ holds for $i=1$, the other cases are similar. Let $f$ be a $k$ valued point of $\underline{\Delta}_{1}\smallsetminus\underline{\Delta}_{2}$ and let $n_{0}$ and $n_{\infty}$ be respectively the multiplicities of 00 and $\infty$ as roots of $f\in k[X,Y]$. Clearly $n_{0}+n_{\infty}\leq 1$ otherwise $f$ would be in $\underline{\Delta}_{2}$.
@@
If $n_{0}+n_{\infty}=0$, then clearly $f$ is not in the image of $p_{1,2}$ and must be in the image of $p_{1,1}$. Write $f=F^{2}G$ where $F\in k[X,Y]_{1}$ and $G\in k[X,Y]_{2}$. Then, since $\mathrm{char}(k)\nmid\deg(G)$, the polynomial $G$ is not a square. Thus $(F,G)$ is uniquely determined. ∎
@@ 3.2 Review of multiplication maps
[…]
@@
$[\Delta_{\mathbb{P}\,W}]=\left[\frac{c(W)}{(1-h_{1})(1-h_{2})}\right]_{\mathrm{rank}(W)-1},$
@@
$\Delta_{\mathbb{P}\,W,*}(h^{k})=h_{1}^{k}\cdot[\Delta_{\mathbb{P}\,W}]$
If $W=E$, then $c(W)=1+c_{1}+c_{2}$, so $[\Delta_{\mathbb{P}\,E}]=h_{1}+h_{2}+c_{1}$ and
@@
$\Delta_{\mathbb{P}\,\mathrm{Sym}^{2}(E),*}(h)$ $=h_{1}^{3}+h_{1}^{2}\otimes h_{2}+h_{1}\otimes h_{2}^{2}+3c_{1}h_{1}\otimes h_{2}+2c_{1}h_{1}^{2}+(2c_{1}^{2}+4c_{2})h_{1}$
@@
The squaring maps are the composition $\mathrm{mult}\circ\Delta_{\mathbb{P}\,W}$, so the result follows from formula (26), after using (23). ∎
@@ 3.3 The classes of the finite subsets
$[X]^{\mathbb{G}_{m}^{2}}=c_{1}^{\mathbb{G}_{m}^{2}}(\mathcal{O}_{\mathbb{P}\,V^{\vee}}(1))-t_{1}$
@@
[$[\{X,Y\}]=\xi_{*}(c_{1}^{\mathbb{G}_{m}^{2}}(\mathcal{O}_{\mathbb{P}\,V^{\vee}}(1))-t_{1})=\xi_{*}(1)c_{1}^{G}(\mathcal{O}_{\mathbb{P}\,V^{\vee}}(1))-\xi_{*}(t_{1})=2s_{1}^{1}-(\beta_{1}+\gamma),$
@@
$0\longrightarrow\mathcal{O}_{BG}\stackrel{{\scriptstyle XY}}{{\longrightarrow}}\mathrm{Sym}^{2}(V_{G}^{\vee})\otimes\Gamma_{G}\otimes\det(V_{G})\longrightarrow Q\longrightarrow 0$ (27)
@@
00{0}${{\mathcal{O}_{\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee})}}}$00{0}${{\mathcal{O}_{\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee})}(-1)\otimes\Gamma_{G}\otimes\det(V_{G})}}$${{\pi^{*}\mathrm{Sym}^{2}(V_{G}^{\vee})\otimes\Gamma_{G}\otimes\det(V_{G})}}$${{\pi^{*}Q}}$00{0}$\scriptstyle{XY}$$\scriptstyle{\psi}$
@@
vanishes. Since the codimension of $\{XY\}\subseteq\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee})$ is $2$, the expected one, we have by [Ful98, Theorem 14.4] that
[$[\{XY\}]=c_{2}(\pi^{*}Q\otimes\mathcal{O}_{\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee})}(1)\otimes\Gamma_{G}^{\vee}\otimes\det(V_{G}^{\vee})).$
@@
[$[\{XY\}]$ $=c_{2}(\pi^{*}Q\otimes\mathcal{O}_{\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee})}(1)\otimes\Gamma_{G}^{\vee}\otimes\det(V_{G}^{\vee}))$
@@
  • [$[\{X^{2},Y^{2}\}]=2s_{2}^{2}-2\beta_{1}s_{2}^{1}+2(\beta_{1}^{2}-2\beta_{2})s_{2}^{0}$.
  • [$[\{X^{2}Y,XY^{2}\}]=2s_{3}^{3}-(3\beta_{1}+\gamma)s_{3}^{2}+2(2\beta_{2}+\beta_{1}^{2})s_{3}^{1}-6\beta_{2}\beta_{1}s_{3}^{0}$.
  • [$[\{X^{3},Y^{3}\}]=2s_{3}^{3}+(\gamma-3\beta_{1})s_{3}^{2}+2(3\beta_{1}^{2}-6\beta_{2})s_{3}^{1}+6\beta_{1}(3\beta_{2}-\beta_{1}^{2})s_{3}^{0}$.
  • [$[\{X^{2}Y^{2}\}]=s_{4}^{4}-2\beta_{1}s_{4}^{3}+2(\beta_{1}^{2}+2\beta_{2})s_{4}^{2}-12\beta_{2}\beta_{1}s_{4}^{1}+24\beta_{2}^{2}s_{4}^{0}$.
  • [$[\{X^{3}Y,XY^{3}\}]=2s_{4}^{4}-4\beta_{1}s_{4}^{3}+6\beta_{1}^{2}s_{4}^{2}-6\beta_{1}^{3}s_{4}^{1}+24\beta_{2}(\beta_{1}^{2}-2\beta_{2})$.
  • [$[\{X^{4},Y^{4}\}]=2s_{4}^{4}-4\beta_{1}s_{4}^{3}+12(\beta_{1}^{2}-2\beta_{2})s_{4}^{2}+24\beta_{1}(\beta_{2}-\beta_{1}^{2})s_{4}^{1}+24(\beta_{1}^{4}+2\beta_{2}^{2}-4\beta_{2}\beta_{1}^{2})s_{4}^{0}$.
@@
In $\mathrm{CH}_{0}(\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee}))$ we have
[$[\{X^{2},Y^{2}\}]$ $=\mathrm{mult}_{*}([\{X,Y\}],[\{X,Y\}])-2[\{XY\}]$
@@
$=2s_{2}^{2}+(-4(\beta_{1}+\gamma)-2(\gamma-\beta_{1}))s_{2}^{1}+\left({2\choose 1}(\beta_{1}+\gamma)^{2}-4\beta_{2}\right)s_{2}^{0}$
@@
$=s_{3}^{3}+(2(\gamma-\beta_{1})-(\beta_{1}+\gamma))s_{3}^{2}+\left(4\beta_{2}-{2\choose 1}(\beta_{1}+\gamma)(\gamma-\beta_{1})\right)s_{3}^{1}-2{3\choose 1}(\beta_{1}+\gamma)\beta_{2}s_{3}^{0}$
@@
$=4s_{3}^{3}-(2(\beta_{1}+\gamma)+4\beta_{1})s_{3}^{2}+\left(2\beta_{1}{2\choose 1}(\beta_{1}+\gamma)+4(\beta_{1}^{2}-2\beta_{2})\right)s_{3}^{1}$
($-{3\choose 1}2(\beta_{1}+\gamma)(\beta_{1}^{2}-2\beta_{2})s_{3}^{0}-2s_{3}^{3}+(3\beta_{1}+\gamma)s_{3}^{2}-2(2\beta_{2}+\beta_{1}^{2})s_{3}^{1}+6\beta_{2}\beta_{1}s_{3}^{0}$
$=2s_{3}^{3}+(\gamma-3\beta_{1})s_{3}^{2}+6(\beta_{1}^{2}-2\beta_{2})s_{3}^{1}+6\beta_{1}(3\beta_{2}-\beta_{1}^{2})s_{3}^{0}$
@@
$=s_{4}^{4}+2(\gamma-\beta_{1})s_{4}^{3}+\left({2\choose 1}(\gamma-\beta_{1})^{2}+4\beta_{2}\right)s_{4}^{2}+4{3\choose 1}\beta_{2}(\gamma-\beta_{1})s_{4}^{1}+4{4\choose 2}\beta_{2}^{2}s_{4}^{0}$
@@
$=2s_{4}^{4}+(2(\gamma-\beta_{1})-2\beta_{1})s_{4}^{3}+\left(4\beta_{2}+2(\beta_{1}^{2}-2\beta_{2})-{2\choose 1}2\beta_{1}(\gamma-\beta_{1})\right)s_{4}^{2}$
+($+{3\choose 1}(-4\beta_{2}\beta_{1}+2(\gamma-\beta_{1})(\beta_{1}^{2}-2\beta_{2}))s_{4}^{1}+{4\choose 2}4(\beta_{1}^{2}-2\beta_{2})\beta_{2}s_{4}^{0}$
$=2s_{4}^{4}-4\beta_{1}s_{4}^{3}+6\beta_{1}^{2}s_{4}^{2}-6\beta_{1}^{3}s_{4}^{1}+24\beta_{2}(\beta_{1}^{2}-2\beta_{2}),$
@@
$=4s_{4}^{4}-8\beta_{1}s_{4}^{3}+\left({2\choose 1}4\beta_{1}^{2}+8(\beta_{1}^{2}-2\beta_{2})\right)s_{4}^{2}-8{3\choose 1}\beta_{1}(\beta_{1}^{2}-2\beta_{2})s_{4}^{1}+4{4\choose 2}(\beta_{1}^{2}-2\beta_{2})^{2}s_{4}^{0}$

### v2

@@ 2.4 Proof of Theorem 2
[…]
@@
We start with showing that $\alpha^{\mathrm{pre}}$ is fully faithful.
@@
The morphism $\alpha^{\mathrm{pre}}$ satisfies condition $(i)$ of Lemma 24.
@@
Since the second map is fully faithful by construction, we only need to prove that the first map is faithful. We will refer to again by the symbol $\alpha^{\mathrm{pre}}$. It is enough to prove that $\alpha^{\mathrm{pre}}(S):\mathcal{X}^{\mathrm{pre}}(S)\to\mathcal{R}_{2}^{\mathrm{st}_{1}}(S)$ is fully faithful for all schemes $S$. Let $u,u^{\prime}:S\to\mathcal{X}^{\mathrm{pre}}$ be objects in $\mathcal{X}$ and let $\alpha^{\mathrm{pre}}(u)=(\pi:C\to S,\eta,\beta)$ and $\alpha^{\mathrm{pre}}(u^{\prime})=(\pi^{\prime}:C^{\prime}\to S,\eta^{\prime},\beta^{\prime})$ be their images under $\alpha^{\mathrm{pre}}$. We wish to show that the induced map
$\mathrm{Mor}_{\mathcal{X}(S)}(u,u^{\prime})\to\mathrm{Mor}_{\mathcal{R}_{2}^{\mathrm{st}_{1}}(S)}((C/S,\eta,\beta),(C^{\prime}/S,\eta^{\prime},\beta^{\prime}))$ (19)
@@
$C{C^{\prime}}$$\mathbb{P}^{1}_{S}$$\mathbb{P}^{1}_{S}$$\scriptstyle{\varphi}$$\scriptstyle{q}$$\scriptstyle{q^{\prime}}$$\scriptstyle{\phi}$ (20)
commutes. The uniqueness of $\phi$ is clear. To show the existence of $\phi$, we consider the corresponding diagram where the vertical arrows are the canonical maps and the lower horizontal arrow is given by the pullback of forms via $\varphi^{*}$
$C{C^{\prime}}$$\mathbb{P}\,(\pi_{*}\omega_{C/S})^{\vee}$$\mathbb{P}\,(\pi^{\prime}_{*}\omega_{C^{\prime}/S})^{\vee}.$$\scriptstyle{\varphi}$$\scriptstyle{p}$$\scriptstyle{p^{\prime}}$$\scriptstyle{\varphi^{*}}$
@@
$C{C^{\prime}}$$\mathbb{P}\,(\pi_{*}\omega_{C/S})^{\vee}$$\mathbb{P}^{1}_{S}$$\mathbb{P}\,(\pi^{\prime}_{*}\omega_{C^{\prime}/S})^{\vee}$$\mathbb{P}^{1}_{S}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{p^{\prime}}$$\scriptstyle{q^{\prime}}$$\scriptstyle{h}$$\scriptstyle{h^{\prime}}$
where $h$ and $h^{\prime}$ are isomorphisms. In these terms, the morphism $\phi$ in diagram (20) is the composition $h^{\prime}\circ\varphi^{*}\circ h^{-1}:\mathbb{P}^{1}_{S}\to\mathbb{P}^{1}_{S}$.
@@
    an isomorphism $\varphi$ between the curve $C$ in (9) and $C^{\prime}$;
@@
Let $\bar{s}$ be a geometric point of $S$. We will prove that (a) and (b) holds in an étale neighborhood of $\bar{s}$. By [Vis98], we know that, perhaps after base-change by an étale cover of $S$, there exists $F$ in $\mathcal{X}^{\mathrm{pre}}$ over $S$ such that there is an isomorphism $\varphi:C\to C^{\prime}$ over $S$. By Lemma 3, we have $\eta^{\prime}_{\bar{s}}=q_{\bar{s}}^{\prime*}\mathcal{O}(1)\otimes\mathcal{O}_{C^{\prime}_{\bar{s}}}(-w_{i}-w_{j})$ for exactly one subset $\{w_{i},w_{j}\}$ of cardinality $2$ of the set of Weierstrass points of $C^{\prime}_{\bar{s}}$. Here $q^{\prime}_{\bar{s}}:C^{\prime}_{\bar{s}}\to\mathbb{P}^{1}$ is any degree $2$ map such that $q_{\bar{s}}^{\prime}(\{w_{i},w_{j}\})=\{0,\infty\}\subseteq\mathbb{P}^{1}$. Up to composing the map (10) with an automorphism of $\mathbb{P}^{1}$ fixing $\{0,\infty\}$ (which amounts to changing $F$ – but not $C$), we may assume that
@@
where $L$ is a line bundle pulled back to $C$ from $S$. Zariski-locally on $S$, the line bundle $L$ is trivial. Up to replacing $S$ by an open subset containing $\bar{s}$ we may then assume that we have an isomorphism $\varphi^{*}\eta^{\prime}\xrightarrow{\tau}\eta$. The diagram
@@
Theorem 2 now follows immediately from Lemma 24 and Propositions 25 and 26. ∎
@@ 3 Computation of the Chow ring of $\mathcal{R}_{2}$
where $\underline{\Delta}\subseteq\mathbb{P}\,\mathrm{Sym}^{4}(V^{\vee})$ is the locus of polynomials (up to non-zero scalars) having either a root at 0 or $\infty$, or a double root in any field extension of the base field $k$.
@@
where $h=c_{1}^{G}(\mathcal{O}_{\mathbb{P}\,\mathrm{Sym}^{4}(V^{\vee})}(1))$, and $\beta_{1}$ and $\gamma$ are defined in Notation 17.
@@
$\mathrm{CH}_{*}^{G}(\underline{\Delta})\to\mathrm{CH}_{*}^{G}(\mathbb{P}\,\mathrm{Sym}^{4}(V^{\vee}))\to\mathrm{CH}^{G}_{*}(\mathbb{P}\,\mathrm{Sym}^{4}(V^{\vee})\smallsetminus\underline{\Delta})\to 0$ (22)
@@ 3.1 A $G$-equivariant envelope for $\underline{\Delta}$
    if $K\supseteq k$ is any extension and $x$ is a $K$-valued point of $X_{i}\smallsetminus X_{i+1}$, then there is a unique $K$-valued point $y$ of $X_{i}^{\prime}$ mapping to $x$.
then the map $p=\sqcup p_{i}:\sqcup_{i=1}^{d}X_{i}^{\prime}\to X$ is a $G$-equivariant envelope.
@@
by $(F,G)\mapsto F^{2}G$, where $G\in\{X^{2},Y^{2},XY\}$. We set $p_{3}=p_{3,1}\sqcup p_{3,2}$ and $\underline{\Delta}_{3}\subseteq\underline{\Delta}_{2}$ to be the image of $p_{3}$.
@@
We will check the conditions of Lemma 30. Condition $(i)$ and $(ii)$ are obvious. We will check that condition $(iii)$ holds for $i=1$, the other cases are similar. Let $f\in K[X,Y]$ be a $K$ valued point of $\underline{\Delta}_{1}\smallsetminus\underline{\Delta}_{2}$ and let $n_{0}$ and $n_{\infty}$ be respectively the multiplicities of 0 and $\infty$ as roots of $f\in K[X,Y]$. Clearly $n_{0}+n_{\infty}\leq 1$ otherwise $f$ would be in $\underline{\Delta}_{2}$.
@@
If $n_{0}+n_{\infty}=0$, then clearly $f$ is not in the image of $p_{1,2}$ and must be in the image of $p_{1,1}$. In fact, by definition, $f$ must have a double root in some extension of $K$, and thus also in $K$ as by assumption $\mathrm{char}(k)$ is bigger than the degree of $f$. Write $f=F^{2}G$ where $F\in K[X,Y]_{1}$ and $G\in K[X,Y]_{2}$. Then, since $f\notin\underline{\Delta}_{2}$, the polynomial $G$ is not a square (in any extension of $K$). Thus $(F,G)$ is uniquely determined. ∎
@@ 3.2 Review of multiplication maps
$\mathrm{mult}_{*}\colon$ $\mathrm{CH}_{*}(\mathbb{P}\,\mathrm{Sym}^{a}(E))\times\mathrm{CH}_{*}(\mathbb{P}\,\mathrm{Sym}^{b}(E))$ $\to{}$ $\mathrm{CH}_{*}(\mathbb{P}\,\mathrm{Sym}^{a+b}(E))$ (26)
$(s_{a}^{\alpha},s_{b}^{\beta})$ $\mapsto{}$ $\binom{a+b-\alpha-\beta}{a-\alpha}s_{a+b}^{\alpha+\beta}$
@@
$\Delta_{\mathbb{P}\,W}]=\left[\frac{c(W)}{(1-h_{1})(1-h_{2})}\right]_{\mathrm{rank}(W)-1},$
@@
$\Delta_{\mathbb{P}\,W,*}(h^{k})=h_{1}^{k}\cdot[\Delta_{\mathbb{P}\,W}].$
If $W=E$, then $c(W)=1+c_{1}+c_{2}$, so $\Delta_{\mathbb{P}\,E}]=h_{1}+h_{2}+c_{1}$ and
@@
$\Delta_{\mathbb{P}\,\mathrm{Sym}^{2}(E),*}(h)$ $=h_{1}^{3}+h_{1}^{2}\otimes h_{2}+h_{1}\otimes h_{2}^{2}+3c_{1}h_{1}\otimes h_{2}+3c_{1}h_{1}^{2}+(2c_{1}^{2}+4c_{2})h_{1}$
@@
where again we have used the projective bundle formula. The squaring maps are the composition $\mathrm{mult}\circ\Delta_{\mathbb{P}\,W}$, so the result follows from formula (26), after using (23). ∎
@@ 3.3 The classes of the finite subsets
$X]^{\mathbb{G}_{m}^{2}}=c_{1}^{\mathbb{G}_{m}^{2}}(\mathcal{O}_{\mathbb{P}\,V^{\vee}}(1))-t_{1}$
@@
$\{X,Y\}]=\xi_{*}(c_{1}^{\mathbb{G}_{m}^{2}}(\mathcal{O}_{\mathbb{P}\,V^{\vee}}(1))-t_{1})=\xi_{*}(1)c_{1}^{G}(\mathcal{O}_{\mathbb{P}\,V^{\vee}}(1))-\xi_{*}(t_{1})=2s_{1}^{1}-(\beta_{1}+\gamma),$
@@
$0\longrightarrow\mathcal{O}_{BG}\stackrel{{\scriptstyle XY}}{{\longrightarrow}}\mathrm{Sym}^{2}(V_{G}^{\vee})\otimes\Gamma_{G}\otimes\det(V_{G})\longrightarrow Q\longrightarrow 0.$ (27)
@@
[…]
@@
vanishes. Since the codimension of $\{XY\}\subseteq\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee})$ is $2$, the expected one, we have by [Ful84, Theorem 14.4] that
$\{XY\}]=c_{2}(\pi^{*}Q\otimes\mathcal{O}_{\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee})}(1)\otimes\Gamma_{G}^{\vee}\otimes\det(V_{G}^{\vee})).$
@@
$[\{XY\}]$ $=c_{2}(\pi^{*}Q\otimes\mathcal{O}_{\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee})}(1)\otimes\Gamma_{G}^{\vee}\otimes\det(V_{G}^{\vee}))$
@@
  • $\{X^{2},Y^{2}\}]=2s_{2}^{2}-2\beta_{1}s_{2}^{1}+2(\beta_{1}^{2}-2\beta_{2})s_{2}^{0}$.
  • $\{X^{2}Y,XY^{2}\}]=2s_{3}^{3}-(3\beta_{1}+\gamma)s_{3}^{2}+2(2\beta_{2}+\beta_{1}^{2})s_{3}^{1}-6\beta_{2}\beta_{1}s_{3}^{0}$.
  • $\{X^{3},Y^{3}\}]=2s_{3}^{3}+(\gamma-3\beta_{1})s_{3}^{2}+2(3\beta_{1}^{2}-6\beta_{2})s_{3}^{1}+6\beta_{1}(3\beta_{2}-\beta_{1}^{2})s_{3}^{0}$.
  • $\{X^{2}Y^{2}\}]=s_{4}^{4}-2\beta_{1}s_{4}^{3}+2(\beta_{1}^{2}+2\beta_{2})s_{4}^{2}-12\beta_{2}\beta_{1}s_{4}^{1}+24\beta_{2}^{2}s_{4}^{0}$.
  • $\{X^{3}Y,XY^{3}\}]=2s_{4}^{4}-4\beta_{1}s_{4}^{3}+6\beta_{1}^{2}s_{4}^{2}-6\beta_{1}^{3}s_{4}^{1}+24\beta_{2}(\beta_{1}^{2}-2\beta_{2})$.
  • $\{X^{4},Y^{4}\}]=2s_{4}^{4}-4\beta_{1}s_{4}^{3}+12(\beta_{1}^{2}-2\beta_{2})s_{4}^{2}+24\beta_{1}(\beta_{2}-\beta_{1}^{2})s_{4}^{1}+24(\beta_{1}^{4}+2\beta_{2}^{2}-4\beta_{2}\beta_{1}^{2})s_{4}^{0}$.
@@
In this proof we will repeatedly apply formula (26) to compute the pushforwards along $\mathrm{mult}$. In $\mathrm{CH}_{0}(\mathbb{P}\,\mathrm{Sym}^{2}(V_{G}^{\vee}))$ we have
$[\{X^{2},Y^{2}\}]$ $=\mathrm{mult}_{*}([\{X,Y\}],[\{X,Y\}])-2[\{XY\}]$
@@
$=2s_{2}^{2}+(-4(\beta_{1}+\gamma)-2(\gamma-\beta_{1}))s_{2}^{1}+\left(\binom{2}{1}(\beta_{1}+\gamma)^{2}-4\beta_{2}\right)s_{2}^{0}$
@@
$=s_{3}^{3}+(2(\gamma-\beta_{1})-(\beta_{1}+\gamma))s_{3}^{2}+\left(4\beta_{2}-\binom{2}{1}(\beta_{1}+\gamma)(\gamma-\beta_{1})\right)s_{3}^{1}-2\binom{3}{1}(\beta_{1}+\gamma)\beta_{2}s_{3}^{0}$
@@
$=4s_{3}^{3}-(2(\beta_{1}+\gamma)+4\beta_{1})s_{3}^{2}+\left(2\beta_{1}\binom{2}{1}(\beta_{1}+\gamma)+4(\beta_{1}^{2}-2\beta_{2})\right)s_{3}^{1}$
$-\binom{3}{1}2(\beta_{1}+\gamma)(\beta_{1}^{2}-2\beta_{2})s_{3}^{0}-2s_{3}^{3}+(3\beta_{1}+\gamma)s_{3}^{2}-2(2\beta_{2}+\beta_{1}^{2})s_{3}^{1}+6\beta_{2}\beta_{1}s_{3}^{0}$
$=2s_{3}^{3}+(\gamma-3\beta_{1})s_{3}^{2}+6(\beta_{1}^{2}-2\beta_{2})s_{3}^{1}+6\beta_{1}(3\beta_{2}-\beta_{1}^{2})s_{3}^{0}.$
@@
$=s_{4}^{4}+2(\gamma-\beta_{1})s_{4}^{3}+\left(\binom{2}{1}(\gamma-\beta_{1})^{2}+4\beta_{2}\right)s_{4}^{2}+4\binom{3}{1}\beta_{2}(\gamma-\beta_{1})s_{4}^{1}+4\binom{4}{2}\beta_{2}^{2}s_{4}^{0}$
@@
$=2s_{4}^{4}+(2(\gamma-\beta_{1})-2\beta_{1})s_{4}^{3}+\left(4\beta_{2}+2(\beta_{1}^{2}-2\beta_{2})-\binom{2}{1}2\beta_{1}(\gamma-\beta_{1})\right)s_{4}^{2}$
$+\binom{3}{1}(-4\beta_{2}\beta_{1}+2(\gamma-\beta_{1})(\beta_{1}^{2}-2\beta_{2}))s_{4}^{1}+\binom{4}{2}4(\beta_{1}^{2}-2\beta_{2})\beta_{2}s_{4}^{0}$
$=2s_{4}^{4}-4\beta_{1}s_{4}^{3}+6\beta_{1}^{2}s_{4}^{2}-6\beta_{1}^{3}s_{4}^{1}+24\beta_{2}(\beta_{1}^{2}-2\beta_{2})s_{4}^{0},$
@@
$=4s_{4}^{4}-8\beta_{1}s_{4}^{3}+\left(\binom{2}{1}4\beta_{1}^{2}+8(\beta_{1}^{2}-2\beta_{2})\right)s_{4}^{2}-8\binom{3}{1}\beta_{1}(\beta_{1}^{2}-2\beta_{2})s_{4}^{1}+4\binom{4}{2}(\beta_{1}^{2}-2\beta_{2})^{2}s_{4}^{0}$